OpenStudy (anonymous):

Find center and radius of circle if: x^2+8x+6y=-16-y^2. Please help me understand this! I don't know how to make it into standard form.

OpenStudy (phi):

do you know how to complete the square for x^2 + 8x

OpenStudy (anonymous):

No

OpenStudy (cwrw238):

x^2 + y^2 + 8x + 6y + 16 = 0 coup;été the square for x^ + 8x and y^2 + 6y and convert it into the form (x - a)^2 + (y - b)^2 = r^2

OpenStudy (phi):

this talks about how to complete the square http://www.khanacademy.org/math/algebra/quadtratics/v/completing-perfect-square-trinomials

OpenStudy (anonymous):

ok I'm downloading the video, thanks.

OpenStudy (phi):

you start with x^2+8x+6y=-16-y^2 or x^2+8x+y^2+ 6y=-16 I'll complete the square for the x terms: take 1/2 the 8 from 8x to get 4. square it to get 16 add 16 -16 (that is zero, so it does not change the equation) x^2 + 8x +16 -16 + y^2 +6y = -16 now rewrite x^2 + 8x +16 as (x+4)^2 to get (x+4)^2 -16 + y^2 +6y = -16 we can move the -16 on the left side to the right side (add +16 to both sides) to get (x+4)^2 + y^2 +6y = 0 compete the square for the y's, and you will get the equation of a circle in standard form